A Young's double slit experiment has the screen pl
Answer #1

a)dsinθ = mλ

0.03*10^-3*sin2 = 2*λ   => λ= 13.5*10^-6 m = 13.5μm

b) tanθ = y/D

y=Dsinθ = 2.6*sin2 = 0.0907 m